OpenStudy (anonymous):

Please help will medal Let m = 2. What is the value of 36m – 16? A. 20 B. 40 C. 56 D. 346

OpenStudy (anonymous):

When a variable touches a number, like in the problem 36m, it means to multiply. So since m = 2 it means 36m = 36 * 2

OpenStudy (anonymous):

@flower2468

OpenStudy (anonymous):

When a variable touches a number, like in the problem 36m, it means to multiply. So since m = 2 it means 36m = 36 * 2

OpenStudy (anonymous):

36*2=72-16=56

OpenStudy (kaleb_m95):

36*2-16=56

OpenStudy (anonymous):

I get medal?

OpenStudy (anonymous):

u too

OpenStudy (kaleb_m95):

yep